CH/Agen/94/04

Agenda

Charissa Group Meeting

To be held on: Friday, 27 May 1994, 9.30 am

To be held at:  Daresbury

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

Charissa “Physics” Meeting

The aim is to review the present state of the analysis of the various experiments and discuss what needs to be done to complete each. We will also hear of the present state of the theoretical developments.

Each speaker should aim to talk for 10 minutes, allowing 10 minutes for discussion.

0930 Introduction BRF
0940 Texas Data BRF
1000 Higher Excitation region BRF
1020 23Na data DLW
1040 LBL 16O+8Be JP
1100 Coffee
1120 16O target CDJ
1140 High Resolution measurement NC
1200 Band termination SPF
1220 Spin at low energy CDJ
1240 Angular Distn SPF
1300 Lunch
1400 12C on 24Mg SMS
1420 40Ca SMS
1440 28Si/32S NC
1500 7 alpha PS
1520 8Be+8Be MF
1540 O3+ SPC
1600 Status of Theory - where next WDMR
1630 Summary and discussion

W.N. Catford (Secretary) 26 May 1994